Angelo's father is 30 years old. This is 10 years older than twice Angelo's age. How old is Angelo?
Found 2 solutions by mananth, stanbon:
Answer by mananth(16946)

Angelo's age =x
2x+10=30
-10
2x=20
/2
x=10 ====> angelo's age

Answer by stanbon(75887)

Angelo's father is 30 years old. This is 10 years older than twice Angelo's age. How old is Angelo?
-----
f = 30 now
---
Equation:
f = 2a+10
Substitute to solve for "a":
30 = 2a+10
2a = 20
a = 10 yrs. old (Angelo's age now)
======================
Cheers,
Stan H.
==========